# Beyond the Permutation Symmetry of Transformers: The Role of Rotation for Model Fusion

### v1

@@ 7 Conclusion
In this paper, we introduced rotation symmetry as a novel type of parameter space symmetry for transformers, extending the concept of symmetry to continuous spaces. Building on this foundation, we proposed a theoretically optimal parameter matching algorithm to enhance the fusion of transformer models in a plug-and-play manner. To validate our approach, we conducted extensive experiments on real-world NLP and vision benchmarks. The results demonstrated that incorporating rotation symmetry effectively and efficiently facilitates transformer model fusion, showcasing our method’s practical utility.
@@ Appendix A Proof
From the proof, we can observe that the optimization problem in Theorem 4.1 has infinite pairs of solutions $\{\bm{R}_{1},\bm{R}_{2}\}$ regarding the value of $\bm{R}_{2}$. In this paper, we let $\bm{R}_{2}=\bm{I}$ for simplicity, which makes model 2 an anchor model. However, the value of $\bm{R}_{2}$ can be further adjusted to benefit some data-dependent model fusion techniques (Singh & Jaggi, 2020; Jin et al., 2023). We leave this part in future works.
Algorithm 1 Matching two self-attention layers.
Appendix B Algorithm Pseudo Code
We provide the pseudo code of the parameter matching algorithm for self-attention layers in Algorithm 1.
Appendix C Comparison with Related Work
Git-Rebasin (Ainsworth et al., 2023) introduces three model fusion algorithms: weight matching, activation matching, and straight-through matching, based on permutation symmetry. While these methods are effective for CNNs and MLPs, they do not easily extend to transformers due to the discrete nature of permutation symmetry. In a discrete space, the parameter matching problem is equivalent to solving a sum of bilinear assignment problems, which is NP-hard (Ainsworth et al., 2023). To address this limitation, we introduce rotation symmetry, which extends the symmetry space to a continuous domain, allowing for a closed-form solution to parameter matching. Unlike permutation symmetry, rotation symmetry provides a more flexible and efficient solution for aligning parameters in transformers.
@@ Appendix C Comparison with Related Work
Appendix D Experimental Details
@@
To address RQ1, we selected three merging methods as baselines: Simple (Wortsman et al., 2022), Fisher (Matena & Raffel, 2022), and RegMean (Jin et al., 2023). Meanwhile, to address RQ2, we further compared the performance of our method with two matching method (Singh & Jaggi, 2020; Ainsworth et al., 2023). The code for the baselines are based on (Jin et al., 2023).
Appendix E Limitations
Similar to previous parameter matching algorithms, our proposed method can only handle the end models with exactly the same architecture. Considering that different pre-trained transformers can have different architectures (most the same, but with slightly different modules), we leave merging different pre-trained transformers as future works.

### v2

@@ 7 Conclusion
In this paper, we introduced rotation symmetry as a novel type of parameter space symmetry for transformers, extending the concept of permutation symmetry to continuous spaces. Building on this foundation, we proposed a theoretically optimal parameter matching algorithm to enhance the fusion of transformer models in a plug-and-play manner. To validate our approach, we conducted extensive experiments on real-world NLP and vision benchmarks. The results demonstrated that incorporating rotation symmetry effectively and efficiently facilitates transformer model fusion, showcasing our method’s practical utility.
@@
Acknowledgment
This work is supported in part by the National Science Foundation (NSF) under grants IIS-2006844, IIS-2144209, IIS-2223769, CNS-2154962, BCS-2228534, and CMMI-2411248; the Office of Naval Research (ONR) under grant N000142412636; and the Commonwealth Cyber Initiative (CCI) under grant VV-1Q24-011.
@@ Appendix A Proof
From the proof, we can observe that the optimization problem in Theorem 4.1 has infinite pairs of solutions $\{\bm{R}_{1},\bm{R}_{2}\}$ regarding the value of $\bm{R}_{2}$. In this paper, we let $\bm{R}_{2}=\bm{I}$ for simplicity, which makes model 2 an anchor model. However, the value of $\bm{R}_{2}$ can be further adjusted to benefit some data-dependent model fusion techniques (Singh & Jaggi, 2020; Jin et al., 2023). We leave this part to future work.
Appendix B Comparison with Related Work
Git Re-Basin (Ainsworth et al., 2023) introduces three model fusion algorithms: weight matching, activation matching, and straight-through matching, based on permutation symmetry. While these methods are effective for CNNs and MLPs, they do not easily extend to transformers due to the discrete nature of permutation symmetry. In a discrete space, the parameter matching problem is equivalent to solving a sum of bilinear assignment problems, which is NP-hard (Ainsworth et al., 2023). To address this limitation, we introduce rotation symmetry, which extends the symmetry space to a continuous domain, allowing for a closed-form solution to parameter matching. Unlike permutation symmetry, rotation symmetry provides a more flexible and efficient solution for aligning parameters in transformers.
@@ Appendix B Comparison with Related Work
Appendix C Experimental Details
@@
We compared the performance of our method with two matching baselines:
  • OTFusion (Singh & Jaggi, 2020) mitigates the positional mismatch of functionally analogous neurons across end models by employing optimal transport to align neurons layer-wise, thus enabling more effective model merging.
  • Git Re-Basin (Ainsworth et al., 2023) matches the parameters of end models through learned permutation symmetries in the parameter space, aiming to transform their parameters into a shared basin within the loss landscape and facilitate seamless merging.
Meanwhile, to address RQ1, we further select six merging methods as baselines:
  • Simple (Wortsman et al., 2022) directly averages the weight matrices of independently pretrained end models.
  • Fisher (Matena & Raffel, 2022) approximates each end model’s posterior as a Gaussian distribution with a precision matrix derived from its Fisher information, thereby casting model merging as a Maximum Likelihood Estimation (MLE) problem over these posterior distributions.
  • RegMean (Jin et al., 2023) formulates model merging as a regression problem, aiming to minimize the Euclidean distance between the predictions of the merged model and individual end models over a set of input samples.
  • OT-WTS (Imfeld et al., 2024) extends OTFusion by introducing soft neuron alignment, which replaces the rigid one-to-one neuron matching with a more flexible, weighted correspondence between neurons. Subsequent model merging is then guided by these soft alignment weights.
  • OT-ACTS and OT-ACTS-EMD (Imfeld et al., 2024) are two variants of OT-WTS that match neurons based on their activations rather than directly aligning parameter matrices. Specifically, OT-ACTS measures activation similarity using Euclidean distance, while OT-ACTS-EMD employs Earth Mover’s Distance for a more fine-grained comparison.
The code implementation is based on (Jin et al., 2023) [^2] https://github.com/bloomberg/dataless-model-merging for NLP tasks and (Imfeld et al., 2024) [^3] https://github.com/graldij/transformer-fusion for vision tasks.
Appendix D Limitations
In this study, we proposed a theoretically optimal parameter matching algorithm based on rotation symmetry. However, the optimality guarantees of our method are subject to several important prerequisites.
First, our theoretical analysis in Theorem 4.1 focuses on the binary model matching case, where the optimal solution is independent of the choice of anchor model. This property, however, does not naturally extend to the multi-model setting. When matching multiple models, a straightforward extension involves selecting a single model as an anchor and aligning the others to it using pairwise rotation-based matching. Yet, in this case, the global optimality depends on the definition of the overall distance metric and may not be preserved. An alternative is iterative pairwise matching, but this introduces path dependency, i.e., the final result can vary depending on the order of merging. Extending our framework to the optimal multi-model matching problem remains an open challenge for future work.
Second, the scope of our parameter matching algorithm is within the rotation symmetry, as defined in Equation 5. While this serves as a tractable and well-motivated instance of parameter space symmetry, we recognize that Transformers exhibit a richer set of structural symmetries beyond permutation, rotation, and rescaling. For example, recent work suggests that geometric properties of normalization and softmax layers may enlarge the symmetry set of previous model layers (Kunin et al., 2021; Lim et al., 2024a). Characterizing the complete symmetry group of transformers remains an open question.
We also considered the possibility of extending rotation matrices to invertible matrices in terms of parameter space symmetries. Although invertible matrices expand the symmetry set, they are less suitable for practical parameter matching. Specifically, our closed-form solution in Equation 9 relies on the orthogonality of the rotation matrices. If replacing them with general invertible matrices, it becomes difficult to derive a closed-form solution for Equation 9, leading to suboptimal solution and increased complexity. Additionally, parameter matching requires computation of the inverse matrix of every $\bm{R}$, which is impractical for invertible matrices. In contrast, orthogonal matrices benefit from computationally efficient inverses (via transposition) and theoretical simplicity.
Finally, we apply the rescaling symmetry after optimizing for rotation-based matching. However, our sequential approach (first optimizing rotation $\bm{R}$, then rescaling weight $\alpha$) does not guarantee global optimality for the joint optimization over $(\bm{R},\alpha)$. Deriving a globally optimal solution would require solving Equation 12 in closed form, which remains non-trivial. Hence, we consider our method as a practical approximation to the joint optimization solution which balances theoretical soundness with computational efficiency. Exploring a fully joint optimization framework for rotation and rescaling symmetries can be a promising direction for future research.